Problem 2 A charge Q is transferred from an initially uncharged plastic ball to an identical ball 27.00 cm away. The force of attraction is then 30.56 mN. How many electrons were transferred from one ball to 1.04E12 Hint: If a charge Q is removed from a neutral sphere, what is the resultant charge? Submit Answer Incorrect. Tries 2/6 Previous Tries Due Friday January 25 11:59 am (ESTHELP PLS!!
Answer #1

According to coloumbs law electric force between two charges is equal to

F = k q1q2 /r2

as charges are having same manitude q1=q2=q

F = kq2 / r2

q = r √(F/K)

q = 27*10-2 √(30.56*10-3 / 9*109)

q=4.975*10-7 Cwe know that q=ne

n= q/e = 4.975*10-7 / (1.6*10^-19) = 3.109*1012 electrons